OpenStudy (anonymous):

help mee ! solve for x , write both solutions. 3x^2+2x-8=0

OpenStudy (anonymous):

looks like someone is attempting to help...

OpenStudy (anonymous):

i need your help. anywasss. plss help. ihts for a semester test.

OpenStudy (anonymous):

a simple way to solve this equation is using the equation to find roots of quadratic equation: \[x = [-b \pm \sqrt{b^{2} - 4ac}]/2a\] other way of solving it would be factoring the equation to: (x + 2)(3x - 4) = 0

OpenStudy (anonymous):

im confuseed noww

OpenStudy (anonymous):

Really, why?

OpenStudy (anonymous):

okay, if you foil the second equation, you should get the same equation as one you put on the question. Just solve (x+2) = 0 and (3x-4) = 0 then solve for x

OpenStudy (anonymous):

thats the point i dont really get what x will be . nothing clickss. im stupid in math sorrry! worst subject since middle school.

OpenStudy (anonymous):

heroo, what can x bee?

OpenStudy (anonymous):

it would be -2 and (4/3)

OpenStudy (anonymous):

Hold on a minute...

OpenStudy (anonymous):

since this is a quadratic equation, there will be two answers. Do you understand the process?

OpenStudy (anonymous):

I think she needs a more elaborate explanation

OpenStudy (anonymous):

to understand the process

OpenStudy (anonymous):

yess sortaa, i really dont gert how you got 4/3

OpenStudy (anonymous):

okay, so you do understand that there are two roots/solutions to x, right?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

okay - so, what i did was to re-write the equation into two smaller equations. and determine solutions based on that smaller equations

OpenStudy (anonymous):

another way of solving this equation is to use the first equation i gave you based on that quadratic equation is expressed in terms of: Ax^2 + Bx + C = 0

OpenStudy (anonymous):

ok thanks , my friend got ihtt already(: